# Parallel transport in rotating frames and projective holonomic quantum computation

### v2

@@ IV   Summary
An oftentimes overlooked fact in nonadiabatic holonomic quantum computation is that the parallel transport condition must be covariantly modified to preserve its physical significance when transitioning between different reference frames. In Sec. II, we demonstrated how to properly modify the parallel transport condition when shifting from the laboratory frame to a rotating reference frame.
@@
In Sec. III, we expanded the gauge theory of nonadiabatic holonomic quantum computation into a projective gauge theory, enabling holonomic quantum computation with projective gates. This new framework eliminates the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. Additionally, we extended the isoholonomic inequality to projective gates and used this extension to derive a sharp estimate for the execution time of projective holonomic gates. This estimate provides a valuable benchmark for developing faster protocols for holonomic quantum gates.

### v3

@@ V   Summary
An oftentimes overlooked fact in nonadiabatic holonomic quantum computation is that the parallel transport condition must be covariantly modified to preserve its physical significance when transitioning between different reference frames. In Section II, we demonstrated how to properly modify the parallel transport condition when shifting from the laboratory frame to a rotating reference frame.
@@
In Section III, we extended the gauge theory of nonadiabatic holonomic quantum computation to a projective gauge theory, enabling holonomic quantum computation with projective gates. This new framework eliminates the problem of global dynamical phases inherent in conventional nonadiabatic holonomic quantum computation. In addition, we extended the isoholonomic inequality to projective gates and used this extension to derive a sharp estimate for the execution time of projective holonomic gates. This estimate provides a valuable benchmark for developing faster protocols for holonomic quantum gates.
Finally, Section IV reviewed the proof in [22] that if the codimension of the computational space is at least as large as its dimension, then any gate can be tightly implemented using a parallel transporting Hamiltonian. By a tight implementation we mean an implementation that saturates the isoholonomic inequality and, hence, the associated execution time estimate. The proof is constructive and provides guidance in designing schemes for tight implementations of holonomic gates.